# Transfer-based Adversarial Poisoning Attacks for Online (MIMO-)Deep Receviers

### v2

@@ III-C Data Poisoning Attacks
The optimization goal of the data poisoning attack is to poison the target with poisoned training data to degrade its test performance, which is shown in Fig. 5(b). Similar to the production of adversarial samples, in the production of poisoning samples, the perturbation $\delta$ is applied to each sample in the training set at $p$-norm and does not exceed the upper bound $\epsilon>0$. Specifically, the test set $\mathcal{V}$ and the training set $\mathcal{T}$ are defined, as well as the poisoned training set $\mathcal{P}(\mathcal{T})$ after applying the perturbation to the data set. Furthermore, $\theta^{*}$ denotes the optimal poisoning parameter of the target classifier with respect to the parameter $\theta$. Thus, the poisoning attack can be modelled as a dual optimization problem as follows:
@@ III-C1 Adversarial Samples as Poisoning Attacks
As previously stated in Section III-B and Section III-C, although both adversarial sample construction for evasion attacks and data poisoning attack sample construction can be attributed to the gradient-based optimization framework, the goals achieved by constructing perturbations for these two types of attacks are different.
Recently, however, researchers have discovered that adversarial samples are also highly effective for poisoning DNNs, a phenomenon known as Adversarial Poisoning [44],[45]. In this case, the poisoning attack optimization problem (5) can also be uniformly expressed in the form of the adversarial sample attack optimization problem (4) in Section III-B. [44] provides a method for creating adversarial poisoning samples to obtain optimal poisoning results. Additionally, [46] shows that antipoisoning attacks can also cause serious harm to the meta-learner in a white-box attack setting. Compared with the dual optimization process of the poisoning attack method in (5), using adversarial samples as poisoning attack samples is more convenient and practically feasible. This is also the basis for the poisoning attack method proposed in this paper.
@@ III-D1 Why Can Attacks Be Transferred?
The concept of attack transferability means that the attack generated for one model can be used against another in a black-box attacks scenario. This phenomenon has been observed and demonstrated in various studies, as referenced in [47]. Given that the approach outlined in this paper can be classified as a black-box attack, it is essential that the attack samples demonstrate the capacity for transferability.
[24] presented an upper bound on the loss theory that arises when black-box transfer occurs. Define $f_{\varphi}$ as the surrogate model with parameter $\varphi$, $f_{\theta}$ as the target model with parameter $\theta$, and $\mathcal{L}(\mathbf{y},\mathbf{s},\varphi)$ as the loss of the input $\mathbf{y}$ in $f_{\varphi}$ against the label $\mathbf{s}$. In consideration of the transferability of evasion attacks (poisoning attacks also take the same form), the optimal adversarial sample, denoted as $\mathbf{y}^{*}$ is obtained by solving (4) on $f_{\varphi}$, with the corresponding optimal perturbation, denoted as $\delta^{*}$. To illustrate, consider the sphere space with $p=2$ and radius denoted as $\epsilon$. The optimal adversarial perturbation obtained on $f_{\varphi}$ can be expressed in (6) as follows:
@@
The left-hand side of the inequality in (7) represents the loss in the black-box attack scenario, while the right-hand side represents the loss in the white-box attack scenario. In the white-box attack scenario, i.e., $f_{\varphi}=f_{\theta}$, inequalities in (7) becomes an equality. The attack achieves its upper bound and has optimal attack effect. Therefore, the effectiveness of the attack when the attack sample is transferred from the surrogate model to the target model is influenced by two factors: The intrinsic adversarial vulnerability of the target model (right-hand side of the inequality in (7)) and the complexity of the surrogate model used to optimize the attack (left side of the inequality in (7)).
The right-hand side of the inequality in (7) shows that a more vulnerable target model has a larger upper bound on the loss, represented by $\epsilon\left|\left|\nabla_{\mathbf{y}}\mathcal{L}(\mathbf{y},\mathbf{s},\theta)\right|\right|_{2}$. The intrinsic complexity of the model measures the learning algorithm’s ability to fit the training data. More complex models, like those without regularization or those prone to overfitting, have more complex parameter spaces and rugged loss landscapes, making them sensitive to input perturbations and susceptible to attacks. For robust models with smaller upper loss bounds, a successful attack requires a higher perturbation limit, reducing the likelihood of bypassing the system’s monitoring. This demonstrates the impact of the intrinsic adversarial vulnerability of the target model on transferability of attacks.
The complexity of the surrogate model used to optimize the attack depends on two main factors: The gradient alignment between the surrogate and the target, and the variance magnitude of the surrogate model’s loss function. These factors are particularly relevant for the left-hand side of the inequality in (7). When the surrogate has better gradient alignment with the target, such as the gradient cosine similarity used in [24], attack samples from the surrogate model exhibit better transferability. This is reflected in $\nabla_{\mathbf{y}}\mathcal{L}(\mathbf{y},\mathbf{s},\varphi)^{\top}\nabla_{\mathbf{y}}\mathcal{L}(\mathbf{y},\mathbf{s},\theta)$, as shown on the left-hand side of the inequality in (7). Additionally, a surrogate model with low variance leads to a more stable optimization process, producing attack samples effective across different target models. In contrast, a large variance leads to an unstable optimization process, leading to attack samples may not match the target model, resulting in failure. Intuitively, on the left-hand side of the inequality in (7), high variance of the loss function increases the corresponding denominator term $\left|\left|\nabla_{\mathbf{y}}\mathcal{L}(\mathbf{y},\mathbf{s},\varphi)\right|\right|_{2}$, which results in smaller upper bounds on the achievability of the transfer.
@@ III-D2 Related Work of Transfer-based Attacks
In summary, when designing transfer-based methods in black-box attacks scenario, one can approach from two perspectives: The inherent adversarial vulnerability of the target model and the complexity of optimizing the surrogate model.
@@
    For the former, certain assumptions about the target model are typically necessary to model the attack objective. For example, one might assume the existence of unstable common vulnerabilities in the integrated model as discussed in [41]. In this scenario, more robust constraints can be applied to the optimization target, or an optimization algorithm with strong generalization ability [41],[43] can be used to optimize the attack samples, making the transfer attack more effective. However, the intrinsic adversarial vulnerability of unknown target is often difficult to identified directly.
@@
    For the latter, typical approaches include using integration-based surrogate model [47], self-integration method based on diverse regularizations [39], and data augmentation [39, 38]. [40] mentioned that alternating different training paradigms (e.g., unsupervised and self-supervised models) as surrogate models when generating poisoning attack samples can improve the transferability. These methods aim to develop more generalized and robust surrogate models, which are then optimized to obtain more effective and transferable attack samples.
@@ IV A poisoning attack framework
As mentioned earlier, in the scenarios discussed in this paper, legitimate deep receivers adapt to fast-varying wireless channels by utilizing pilot and updating model parameters using online learning methods. However, this online update mechanism, aimed at adapting to local channel variations, faces the threat of sample input poisoning.
In the attack framework of the malicious user, the legitimate online deep receiver becomes the target model, which is updated online to adapt to local variations in the channel for better local performance. However, this process is inherently not robust and the local adaptation can result in overfitting [48]. From this point onwards, the attack target of this paper is similar to [49], which causes catastrophic forgetting of the target by constructing poisoned samples. According to the analysis of attack transferability in Section III-D, this overfitting is the source of the target model’s intrinsic adversarial vulnerability. This implies that a malicious user can effectively attack the target model by optimizing a surrogate model and generating corresponding adversarial perturbations.
Therefore, this section proposes a poisoning attack framework targeting online deep receivers. The core idea is to poison the model training and updating phases, resulting in a poisoned model after a certain period, leading to performance degradation. Specifically, based on the behavioral patterns of malicious users described in Section II-C, the poisoning attack framework and the method for generating the poisoning attack samples can be refined into below three steps, as illustrated in Fig. 7.
@@ 1. Step 1:
    The malicious user collects communication data (e.g., pilot) from the wireless channel and produces a joint learning dataset. This dataset is used to train a surrogate model for attacking the target (the legitimate deep receiver).
@@ IV-A Surrogate Model Selection
As discussed earlier in Section III-D regarding attack transferability, black-box attacks require the malicious user to consider the degree of gradient alignment between the surrogate model and the target model to ensure the attack’s generalization. To achieve this, it is important to avoid using surrogate models that are overly specialized for specific problems. This strategy ensures consistent compatibility with various types of deep receivers, thereby enhancing the attack’s overall effectiveness. Consequently, this paper adopts a generic DNN architecture, e.g., feedforward neural networks, as the surrogate model for the attack, as illustrated in Step 1 of Fig. 6.
@@ IV-B Joint Learning for Training of Surrogate Models
From the perspective of a malicious user, it is essential to select a suitable surrogate model architecture similar to the target model and to effectively train and optimize the surrogate model. In the proposed attack framework, joint learning is used to train the surrogate model. This approach utilizes data collected under various channel conditions to train the DNN, enabling it to adapt to dynamic channels [13, 14, 15].
It is important to note that, unlike legitimate communicating parties that have to use online learning methods to adapt to time-varying wireless channels, the malicious user has preparation time to eavesdrop communications and collect large samples to train a robust model. Furthermore, the use of joint learning, instead of channel state information as input or online learning [15], addresses the issue of attack generality at the data level, avoiding over-specialized surrogate models in black-box attack transfers, as previously discussed in Section IV-A. Finally, in terms of training effectiveness, joint learning meets the data volume requirements of deep learning, making the surrogate model more robust compared to the target model using online learning. This results in attack patterns with more stable and effective attack results. Conversely, the robust learning process of the target model mitigates the impact of poisoning. To demonstrate this, we will adjust the pilot size in Section V-E.
Specifically, as illustrated in Step 2 of Fig. 6, the malicious user deploys joint learning to train a DNN (i.e., the surrogate model) utilizing a substantial quantity of communication data (e.g., pilot) amassed from authorised transceivers. The objective is for this DNN to learn a mapping that is applicable to the majority of channel states, reflecting the input-to-output mapping relationship of the authorised deep receivers (i.e., the target model). The data that is jointly learned comprises two categories: Communication data under disparate channel distribution conditions and communication data under varying signal-to-noise ratio (SNR) conditions within the same channel. Finally, to produce effective adversarial poisoning samples, malicious user need to consider suitable attack generation methods based on the surrogate model, as detailed in Section IV-C.
@@ V NUMERICAL EVALUATIONS
In this section, we conduct a numerical evaluation of the proposed poisoning attack method aimed at disrupting the online adaptation process of the deep receiver. We first provide provides an overview of the parameter settings employed in the simulation experiments in Sections V-A to V-D. These parameters include the channel models, the deep receivers, the online training methods and the poisoning attack method, which are all evaluated in this paper. Subsequently, the simulation results for our evaluated deep receivers are presented under the following conditions: A linear time-varying synthetic channel, a nonlinear time-varying synthetic channel, a linear static synthetic channel, and the time-varying COST 2100 channel (Section V-E). Finally, the experimental results of the proposed attack method under the four channel settings are summarised and discussed in Section V-F.
@@ V-A Evaluated Channel Models
The deep receiver operates on a discrete memoryless MIMO channel. The number of transmitting and receiving antennas is set to $N_{tx}=N_{rx}=4$. The experimental evaluation channel model comprises synthetic channels [20] and COST 2100 channels [50]. Fig. 7 illustrates the four channel tapping coefficients for a randomly selected user in a multiuser system over 100 blocks of data transmission.
@@
In the context of linear channels, the input-output relationship is expressed by the (1) given in Section II-A. In the context of a nonlinear channel model [20], the input-output relationship is represented by
@@ V-B Evaluated Deep Receivers
    DeepSIC: SIC estimates and cancels the interfering signals at each iteration to revise previous estimates, while DeepSIC unfolds the SIC iterative process and replaces it with a sub-neural network to improve the performance of each iteration. The function of each sub-neural network is to enhance the reliability of the present estimation by utilizing the received symbols and the preceding iteration’s output confidence. This design enables DeepSIC to attain high reliability even with a limited number of training data[1, 20]. Furthermore, DeepSIC presented in this paper incorporates 3 iterations, resulting in a total of $3\times N_{tx}$ sub-networks. In this configuration, each sub-network is a two-layer fully connected layer network. The first layer has a dimension of $(N_{rx}+N_{tx}-1)\times 64$, and the second layer has a dimension of $64\times|S|$, where $|S|$ is the size of the set of symbols to be transmitted. For example, $|S|=4$ when using QPSK transmission. The activation function employed in the initial layer of each subnetwork is ReLU, while the second layer utilizes softmax classification.
@@
    Black-box DNN Detector: The black-box DNN architecture comprises four fully-connected layers and a softmax classification header. The input and output dimensions are $N_{rx}\times 60,60\times 60,60\times 60$ and $60\times|S|^{N_{tx}}$, respectively, for each layer of the network.
@@
    ResNet Detector: The ResNet detector employed in this paper consists of 10 layers of residual blocks, with each block comprising two convolutional layers that use 3$\times$3 kernels, one-pixel padding on both sides, and no bias terms, with a ReLU activation function in between. Each convolutional layer is followed by 2D batch normalization.

### v3

@@ III-C Data Poisoning Attacks
The optimization goal of the data poisoning attacks is to poison the target with poisoned training data to degrade its test performance, which is shown in Fig. 5(b). Similar to the generation of adversarial samples, in data poisoning attacks, a perturbation $\delta$ is applied to each sample in the training set under the $p$-norm constraint, ensuring that the perturbation magnitude does not exceed an upper bound $\epsilon$. Specifically, the test set $\mathcal{V}$ and the training set $\mathcal{T}$ are defined, as well as the poisoned training set $\mathcal{P}(\mathcal{T})$ after applying the perturbation to the data set. Furthermore, $\theta^{*}$ denotes the optimal poisoning parameter of the target classifier with respect to the parameter $\theta$. Thus, the poisoning attack can be modelled as a dual optimization problem as follows:
@@ III-C1 Adversarial Samples as Poisoning Attacks
As previously stated in Section III-B and Section III-C, although both construction of evasion attacks and poisoning attacks can be framed as gradient-based optimization problems, the goals achieved by constructing perturbations for these two types of attacks are different. Recently, however, researchers have discovered that adversarial samples are also highly effective for poisoning DNNs, a phenomenon known as Adversarial Poisoning [44],[45]. In this case, the poisoning attack optimization problem (5) can also be uniformly expressed in the form of the adversarial sample attack optimization problem (4) in Section III-B. [44] provides a method for creating adversarial poisoning samples to obtain optimal poisoning results. Additionally, [46] shows that adversarial poisoning attacks can also cause serious harm to the meta-learner in a white-box attack setting. Compared with the dual optimization process of the poisoning attack method in (5), using adversarial samples as poisoning attack samples is more convenient and practically feasible. This is also the basis for the poisoning attack method proposed in this paper.
@@ III-D1 Why Can Attacks Be Transferred?
Attack transferability means that the attack generated for a model may be equally effective for other models that have not seen that attack before. This phenomenon has been observed and demonstrated in [47]. The proposed attack method is a black-box attack. Therefore, it is crucial to analyze the source of the attack portability.
[24] presented an upper bound on the loss when black-box transfer occurs. Define $f_{\varphi}$ as the surrogate model with parameter $\varphi$, $f_{\theta}$ as the target model with parameter $\theta$, and $\mathcal{L}(\mathbf{y},\mathbf{s},\varphi)$ as the loss of the input $\mathbf{y}$ in $f_{\varphi}$ against the label $\mathbf{s}$. In consideration of the transferability of evasion attacks (poisoning attacks also take the same form), the optimal adversarial sample, denoted as $\mathbf{y}^{*}$ is obtained by solving (4) on $f_{\varphi}$, with the corresponding optimal perturbation, denoted as $\delta^{*}$. To illustrate, consider the sphere space with $p=2$ and radius denoted as $\epsilon$. The optimal adversarial perturbation obtained on $f_{\varphi}$ can be expressed in (6) as follows:
@@
The left-hand side of the inequality in (7) represents the loss in the black-box attack scenario, while the right-hand side represents the loss in the white-box attack scenario. In the white-box attack scenario, i.e., when $f_{\varphi}=f_{\theta}$, the inequalities in (7) becomes an equality. At this point, the attack achieves its upper bound and has optimal effectiveness.
Thus, the effectiveness of transferring an attack sample from the surrogate model to the target model is influenced by two factors: The intrinsic adversarial vulnerability of the target model (right-hand side of the inequality in (7)) and the complexity of the surrogate model used to optimize the attack (left-hand side of the inequality in (7)). The right-hand side of the inequality in (7) shows that a more vulnerable target model has a larger upper bound on the loss, represented by $\epsilon\left|\left|\nabla_{\mathbf{y}}\mathcal{L}(\mathbf{y},\mathbf{s},\theta)\right|\right|_{2}$. Here, the intrinsic complexity of the model measures the learning algorithm’s ability to fit the training data. More complex models, like those without regularization or those prone to overfitting, have more intricate parameter spaces and rugged loss landscapes, making them sensitive to input perturbations. For robust models with smaller upper loss bounds, a successful attack requires a higher perturbation limit, reducing the likelihood of bypassing the system’s monitoring. This demonstrates the impact of the intrinsic adversarial vulnerability of the target model on transferability of attacks.
The complexity of the surrogate model used to optimize the attack depends on two main factors: The gradient alignment between the surrogate and the target, and the variance magnitude of the surrogate model’s loss function. These factors are particularly relevant for the left-hand side of the inequality in (7). When the surrogate has better gradient alignment with the target, such as higher gradient cosine similarity [24], the attack samples from the surrogate model exhibit better transferability. This is reflected in $\nabla_{\mathbf{y}}\mathcal{L}(\mathbf{y},\mathbf{s},\varphi)^{\top}\nabla_{\mathbf{y}}\mathcal{L}(\mathbf{y},\mathbf{s},\theta)$, as shown on the left-hand side of the inequality in (7). Additionally, a surrogate model with low variance leads to a more stable optimization process, producing attack samples effective across different target models. In contrast, a large variance leads to an unstable optimization process, leading to attack samples may not match the target model, resulting in failure. Intuitively, on the left-hand side of the inequality in (7), high variance of the loss function increases the corresponding denominator term $\left|\left|\nabla_{\mathbf{y}}\mathcal{L}(\mathbf{y},\mathbf{s},\varphi)\right|\right|_{2}$, which results in smaller upper bounds on the achievability of the transfer.
@@ III-D2 Related Work of Transfer-based Attacks
In summary, transfer-based attacks can be approached from two perspectives: The inherent adversarial vulnerability of the target model and the complexity of optimizing the surrogate model.
@@
    For the former, certain assumptions about the target model are typically necessary to model the attack objective. For example, [41] assumed unstable common weakness in the model ensemble. In this scenario, stronger constraints can be applied to the optimization target, such as sharpness-aware minimization[41] or momentum methods [43], to make the transfer attack more effective. However, the intrinsic adversarial vulnerability of unknown target is often difficult to identified directly.
@@
    For the latter, approaches include using ensembled surrogate model [47], self-ensembled method [39], and data augmentation [39, 38]. [40] suggested alternating different training paradigms (e.g., unsupervised and self-supervised) to enhance transferability in poisoning attacks. These methods aim to build more generalized and robust surrogate models, which are then optimized to obtain more effective and transferable attack samples.
@@ IV A poisoning attack framework
As mentioned earlier, in the scenarios discussed in this paper, legitimate deep receivers adapt to fast-varying wireless channels by utilizing pilot and updating model parameters through online learning. However, this online update mechanism, designed to adapt to local channel variations, is vulnerable to sample input poisoning.
In the attack framework of the malicious user, the legitimate online deep receiver becomes the target model, which is updated online to adapt to local channel variations for better performance. However, this process is inherently not robust and can result in overfitting [48]. From this point onwards, the attack target of this paper is similar to [49], which causes catastrophic forgetting of the target by constructing poisoned samples. According to the analysis in Section III-D, this overfitting is the source of the target model’s intrinsic adversarial vulnerability. This implies that a malicious user can effectively attack the target model by optimizing a surrogate model and generating corresponding adversarial perturbations.
Therefore, this section proposes a poisoning attack framework targeting online deep receivers. The core idea is to poison the model training and updating phases, resulting in a degraded model after a certain period. Specifically, based on the behavioral patterns of malicious user described in Section II-C, the poisoning attack framework and the method for generating the poisoning attack samples can be summarized in below 3 steps, as illustrated in Fig. 7.
@@ 1. Step 1:
    The malicious user collects communication data (e.g., pilot) from the wireless channel and produces a joint learning dataset. This dataset is used to train a surrogate model for attacking the target (i.e., the legitimate deep receiver).
@@ IV-A Surrogate Model Selection
As discussed in Section III-D, black-box attacks rely on the gradient alignment between the surrogate and target models to ensure generalization. To achieve this, it is crucial to avoid using surrogate models that are too specialized for specific tasks. This strategy improves compatibility with different deep receivers, enhancing the attack’s effectiveness. Therefore, this paper employs a generic DNN architecture, such as feedforward neural networks, as the surrogate model, as shown in Step 1 of Fig. 6.
@@ IV-B Joint Learning for Training of Surrogate Models
From the perspective of a malicious user, it is essential to select a suitable surrogate model architecture similar to the target model and to effectively train and optimize the surrogate model. In the proposed attack framework, joint learning is used to train the surrogate model. This approach utilizes data collected under various channel conditions to train the DNN, enabling it to adapt to dynamic channels [13, 15]. Unlike legitimate communicating parties that have to use online learning methods to adapt to time-varying wireless channels, the malicious user can pre-collect large datasets to train a robust model. Furthermore, the use of joint learning, instead of using channel state information or online learning [15], addresses the issue of attack generality at the data level, avoiding over-specialized surrogate models in black-box attack transfers, as discussed in Section IV-A. Joint learning also satisfies the data volume requirements of deep learning, making the surrogate model more robust than the target model, which uses online learning. This results in attack patterns with more stable and effective attack results. Conversely, the robust learning process of the target model mitigates the impact of poisoning, as demonstrated by pilot size adjustments in Section V-E.
Specifically, as shown in Step 2 of Fig. 6, the malicious user employs joint learning to train a DNN (i.e., the surrogate model) utilizing a large amount of communication data (e.g., pilot) amassed from legitimate transceivers. The DNN learns a mapping applicable to most channel states, reflecting the input-output relationship of the target deep receivers. The training data includes two types: communication data from different channel distributions and data from varying signal-to-noise ratio (SNR) conditions within the same channel. Finally, to generate effective adversarial poisoning samples, suitable attack generation methods must be considered, as detailed in Section IV-C.
@@ V NUMERICAL EVALUATIONS
In this section, we numerically evaluate the proposed poisoning attack aimed at disrupting the deep receiver’s online adaptation process. First, we outline the parameter settings used in the simulations, including channel models, deep receivers, online training methods, and the poisoning attack method (Sections V-A to V-D). Next, simulation results are presented for deep receivers under the following conditions: linear and nonlinear time-varying synthetic channels, a linear static synthetic channel, and the time-varying COST 2100 channel (Section V-E). Lastly, the experimental results across all four channel settings are summarized and discussed in Section V-F.
@@ V-A Evaluated Channel Models
The deep receiver operates on a discrete memoryless MIMO channel. The number of transmitting and receiving antennas is set to $N_{tx}$$=$$N_{rx}$$=$4. The experimental evaluation channel model comprises synthetic channels [20] and COST 2100 channels [50]. Fig. 7 illustrates the four channel tapping coefficients for a randomly selected user in a multiuser system over 100 blocks of data transmission. In the context of linear channels, the input-output relationship is expressed by the (1) given in Section II-A. In the context of a nonlinear channel model [20], the input-output relationship is represented by
@@ V-B Evaluated Deep Receivers
    DeepSIC: DeepSIC unfolds the traditional SIC iterative process and replaces each iteration with a sub-neural network to improve performance. Each sub-neural network enhances the reliability of the current estimation using received symbols and the output confidence from the previous iteration. This design enables DeepSIC to achieve high reliability even with limited training data[1, 20]. In this paper, DeepSIC includes 3 iterations, resulting in a total of $3\times N_{tx}$ sub-networks. Each sub-network is a two-layer fully connected layer network. The first layer has a dimension of $(N_{rx}+N_{tx}-1)\times 64$, and the second layer has a dimension of $64\times|S|$, where $|S|$ is the size of the set of symbols to be transmitted. For example, $|S|=4$ when using QPSK transmission. The activation function employed in the initial layer of each subnetwork is ReLU, while the second layer utilizes softmax classification.
@@
    Black-box DNN Detector: The black-box DNN architecture consists of 4 fully-connected layers and a softmax classification header. The dimensions of the layers are $N_{rx}\times 60,60\times 60,60\times 60$ and $60\times|S|^{N_{tx}}$, respectively.
@@
    ResNet Detector: The ResNet detector employed in this paper comprises 10 layers of residual blocks, each containing 2 convolutional layers with 3$\times$3 kernels, one-pixel padding on both sides, and no bias terms. A ReLU activation function is used between the layers, and each convolutional layer is followed by 2D batch normalization.